OpenStudy (anonymous):

What is the equation of the line parellel to y=1/5x -10 and passes through (15,-16)?

OpenStudy (abb0t):

If the line is parallel, then that means it has the same slope. Plug in the (x,y) coordinate and the slope and solve for b.

OpenStudy (abb0t):

Try it first. You're already given the information: m = 1/5 x = 15 y = -16 y = mx + b b = y-mx